4.10. Let T be a linear transformation on a vector space V satisfying T-T2 = id. Show that T is invertible.
Answer #1

given that T-T²=1 (I-T) T=1 → ST=I where S=I-T - T Isleft Inveible Similarly T-T²=I TCT-1)=I TU I with U=1-T TIs right invert